# RoToR: Towards More Reliable Responses for Order-Invariant Inputs

### v1

@@ Abstract
Mitigating positional bias of language models (LMs) for listwise inputs is a well-known and important problem (e.g., lost-in-the-middle). While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited. In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems. To overcome, we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks. On the Lost in the middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner. [^1] https://github.com/soyoung97/RoToR
@@ 1 Introduction
Language conveys meaning in part through positional information, such as word placement and sentence structure. Given this nature, Language Models (LMs) that learn from human language are trained sensitive to positional information related to the ordering of segments. However, there are some listwise inputs that require neutrality to positional information. For example, for inputs such as sets, tables, databases, or multiple-choice questions, the ordering of the input segments—e.g., rows in a table or elements in an unordered set—require an order-agnostic understanding. We refer to such inputs as “order-invariant inputs,” which LMs reportedly struggle on. For example, in LLM-as-a judge scenarios, they exhibit a preference of up to 75% for the first answer in pairwise inputs Zheng et al. (2024b), and ranking between LMs can change up to 8 positions in different orderings of multiple choice questions on MMLU Alzahrani et al. (2024). Such results question the reliability of LM outputs on order-invariant inputs. Meanwhile, existing methods for enforcing invariance to LMs showed limited effectiveness in real-world tasks, which we hypothesize to arise from the following limitations.
@@
First, training and inference distribution mismatch due to the positional ID re-assignment of zero-shot order-invariant LMs: Fig. 1 illustrates how self-attention is altered in there models. Unlike the original non-invariant model which always assigns position IDs in a causal, ascending manner, order-invariant models either eliminate inter-segment attention, such as PCW Ratner et al. (2023) in Fig. 1a, or re-assign position IDs as in PINE Wang et al. (2024) in Fig. 1b, re-ordering segments using an importance score, placing similar segments closer to the query. For each query segment, it computes segment-wise query-key attention (for each attention head in each decoder layer) and re-assigns position IDs of segments as keys. This query-dependent segment ordering leads to excessively frequent alterations on positional ID assignments. Frequent re-assignments can also confuse the model and risk collisions which violates the invariance property (e.g., multiple segments having the same importance score).
@@
To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment. Ours is named RoToR, inspired by the word rotary to express circular assignment, and also a palindrome, to reflect order invariance. Fig. 1c contrasts with PINE in Fig. 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation. The ordering of segments on suffix tokens remain in a fixed order, since it does not rely on their similarity to the query. Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
@@
Second, in practical listwise inputs, order-invariant task may include order-sensitive inputs which requires order-specific understanding. For example, the (d) None of the above option in MMLU cannot be reordered. For such a “mixed” nature requires to handle each of them adaptively by a simple Selective Routing method. Selective Routing adapts to input by routing between two models, invariant and non-invariant (original), based on their confidence scores of predictions. Experiments on the MMLU benchmark show that Selective Routing effectively handles datasets with order-invariant and sensitive inputs, and achieve better order robustness while maintaining the original performance.
@@ 2.1 Positional bias of LLMs
Considerations for decoder-only LMs. While successful approaches are presented to mitigate this issue for encoder-only Yang et al. (2022) and encoder-decoder Yen et al. (2024); Cai et al. (2023) models, leaving decoder-only models, which account for most of the current LLMs, for more consideration. In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al. (2019), transformer decoders use causal attention in order to learn causal relation signals, which is not invariant by nature Haviv et al. (2022a). Therefore, positional bias for decoder-only models is known to stem from both positional encoding and causal attention mask Yu et al. (2024); Wang et al. (2024).
@@ 3.1 Baseline: Order-invariant causal LMs
In this section, we briefly overview the existing work on turning decoder-only LLM architectures into order-invariant models by adjusting attention mechanism, and review their limitations.
@@
A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism, by letting each segment attend to all other segments. However, to allow this within decoder-only models with causal attention and still achieve order invariance, PINE modifies the multi-head self-attention procedure to create an ‘illusion’: it treats each query segment as if it were the final segment in the input list, so that it can attend to all other segments as keys. For each query segment, the ordering among the key segments is determined by their attention scores (without positional embeddings, $\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that segments with stronger relevance to the query appear closer. Meanwhile, the tokens within a query segment still follow causal ordering. Fig. 2 illustrates this with the input [T1‘‘Given’’, S1[‘‘Apple’’], S2[‘‘Ban’’, ‘‘ana’’], S3[‘‘Orange’’], T6‘‘which one’’, .. T10‘‘red?’’]. Prefix tokens ‘‘Given’’ and suffix tokens ‘‘which one .. red?’’ remain in their original positions and follow normal causal attention, equally attending to all segments. In contrast, the segments in the middle (S1 - S3) require the order-invariant mechanism. PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID. For Case 1: When “ana” (T4) in segment S2 acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments. Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions). Case 2: When “ana” serves as a key, the placement of S2 depends on its $\mathrm{Attn}_{\mathrm{NoPoS}}$ scores with the query segment. If “banana” (S2) is more relevant to the query segment “apple” (S1) than “orange” (S3), (i.e., $\mathbf{Attn_{NoPos}}$(S1, S2) > $\mathbf{Attn_{NoPos}}$ (S1, S3)) S2 is placed nearer to S1, with “ban” and “ana” preserving their internal sequence. For prefix, suffix, and generated tokens, they do not participate in order invariance, so they are always placed at their standard causal positions. However, the same dynamic reordering applies to segment tokens (T2–T5) when computing attention scores of suffix (T6–T10) and the generated token (T11) as query.
@@ 4 Experiment setup
Gold idx at: 0 4 9 avg. 0 4 9 14 19 avg. 0 4 9 14 19 24 29 avg.
Indexing bias present
Indexing bias removed
@@
WebQSP (Count = 1466) Mintaka (Count = 4000)
N=30 N=50 N=30 N=50
Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg.
Initial (no shuffling of segments)
Orig. 71.3 43.2 58.3 57.6 71.6 43.0 58.1 57.5 50.2 44.0 51.9 48.7 50.0 44.0 51.7 48.6
PINE 72.0 44.1 59.0 58.4 72.4 44.4 58.5 58.4 51.5 45.0 52.6 49.7 51.6 45.1 52.6 49.8
RoToR 73.7 43.4 59.2 58.8 73.9 43.5 59.0 58.8 53.1 46.5 54.1 51.2 52.9 46.0 53.6 50.8
RoToR (MonoT5) 72.9 43.6 59.2 58.6 73.9 43.9 59.1 59.0 51.6 45.0 52.5 49.7 52.4 45.4 52.8 50.2
RoToR (Freq) 73.5 43.5 59.3 58.8 74.1 43.9 59.3 59.1 52.6 46.1 53.7 50.8 53.1 46.4 53.7 51.0
@@
Orig. 70.9 42.4 57.8 57.0 $\pm$ 0.5 70.7 41.6 57.2 56.5 $\pm$ 0.2 49.5 43.3 51.0 47.9 $\pm$ 0.0 49.7 43.5 51.0 48.1 $\pm$ 0.4
PINE 71.9 43.9 59.1 58.1 $\pm$ 0.0 72.5 44.6 58.5 58.5 $\pm$ 0.3 51.8 45.2 52.8 50.0 $\pm$ 0.1 51.8 45.3 52.7 49.9 $\pm$ 0.2
RoToR 73.6 43.2 59.1 58.7 $\pm$ 0.1 74.0 43.3 58.8 58.7 $\pm$ 0.1 52.8 46.2 53.8 50.9 $\pm$ 0.0 52.7 45.9 53.5 50.7 $\pm$ 0.0
RoToR (MonoT5) 72.6 43.4 59.2 58.3 $\pm$ 0.1 73.9 43.7 58.9 58.8 $\pm$ 0.1 51.6 45.0 52.6 49.7 $\pm$ 0.1 52.2 45.2 52.8 50.1 $\pm$ 0.2
RoToR (Freq) 72.9 43.2 59.0 58.4 $\pm$ 0.0 73.7 43.8 59.1 58.9 $\pm$ 0.1 52.5 45.9 53.5 50.6 $\pm$ 0.1 53.1 46.4 53.7 51.1 $\pm$ 0.0
Table 2: Knowledge Graph Question Answering (KGQA) results, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
Initial
ordering
Reversed
ordering
Avg. of all possible
(4!-1) re-orderings
Baseline: Single use of order-invariant models
Orig. 68.3 64.8 65.5 $\pm$ 1.0
@@
PINE 64.8 63.3 63.6 $\pm$ 0.7
RoToR 63.2 62.6 62.8 $\pm$ 0.7
RoToR (MonoT5) 64.2 62.9 63.5 $\pm$ 0.5
RoToR (Freq) 64.3 63.5 63.6 $\pm$ 0.6
With Selective Routing
Orig. 68.3 64.8 65.5 $\pm$ 1.0
Orig + PINE 68.5 65.1 65.7 $\pm$ 0.9
Orig + RoToR 68.5 65.0 65.7 $\pm$ 0.9
Orig + RoToR (MonoT5) 68.4 65.2 65.8 $\pm$ 0.9
Orig + RoToR (Freq) 68.5 65.3 65.8 $\pm$ 0.8
Orig + RoToR (Optimal) 75.0 71.9 72.7 $\pm$ 1.0
Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing, which includes both order-invariant and order-sensitive examples. RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optimal) implies the potential for further improvements.

### v2

@@ Abstract
Mitigating positional bias of language models (LMs) for listwise inputs is a well-known and important problem (e.g., lost-in-the-middle). While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited. In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems. Then, to overcome these issues we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks. On the Lost in the middle (LitM), Knowledge Graph QA (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner. [^1] https://github.com/soyoung97/RoToR
@@ 1 Introduction
Language conveys meaning in part through positional information, such as word placement and sentence structure. Given this nature, Language Models (LMs) that learn from human language are trained sensitive to positional information related to the ordering of segments. However, there are some listwise inputs that require neutrality to positional information. For example, for inputs such as sets, tables, databases, or multiple-choice questions, the ordering of the input segments—e.g., rows in a table or elements in an unordered set—require an order-agnostic understanding. We refer to such inputs as “order-invariant inputs,” on which LMs reportedly struggle. For example, in LLM-as-a-judge scenarios, LMs exhibit a preference of up to 75% for the first answer in pairwise inputs Zheng et al. (2024b), and ranking between LMs can change up to 8 positions in different orderings of multiple choice questions on MMLU Alzahrani et al. (2024). Such results question the reliability of LMs on order-invariant inputs. Meanwhile, existing methods for enforcing invariance to LMs showed limited effectiveness in real-world tasks, which we hypothesize to arise from the following limitations.
@@
First, training and inference distribution mismatch due to the positional ID re-assignment of zero-shot order-invariant LMs: Fig. 1 illustrates how self-attention is altered in these models. Unlike the original non-invariant model which always assigns position IDs in a causal, ascending manner, order-invariant models either eliminate inter-segment attention, such as PCW Ratner et al. (2023) in Fig. 1a, or re-assign position IDs as in PINE Wang et al. (2024) in Fig. 1b, re-ordering segments using pairwise similarity, placing similar segments closer to the query. For each query segment, it computes segment-wise query-key attention (for each attention head in each decoder layer) and re-assigns position IDs of segments as keys. This query-dependent segment ordering leads to excessively frequent alterations of positional ID assignments. Frequent re-assignments can also confuse the model and risk collisions which violate the invariance property (e.g., multiple key segments having the same similarity to a query).
@@
To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment. Ours is named RoToR, inspired by the word rotary to express circular assignment, and also a palindrome, to reflect order invariance. Fig. 1c contrasts with PINE in Fig. 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation. The ordering of segments on suffix tokens remains in a fixed order, since it does not rely on their similarity to the query. Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
@@
Second, for practical listwise inputs, order-invariant tasks may partially include order-sensitive inputs that require order-specific understanding. For example, the (d) None of the above option in MMLU cannot be reordered. Such a “mixed” nature requires handling each of the cases adaptively, for which we propose a simple Selective Routing method. Selective Routing adapts to a given input by routing between two models, invariant and non-invariant (original), based on the confidence scores of their predictions. Experiments on the MMLU benchmark show that Selective Routing effectively handles datasets with order-invariant and sensitive inputs, and achieves better order robustness while maintaining the original performance.
@@ 2.1 Positional bias of LLMs
Considerations for decoder-only LMs. While successful approaches are presented to mitigate this issue for encoder-only Yang et al. (2022) and encoder-decoder Yen et al. (2024); Cai et al. (2023) models, they leave decoder-only models, which account for most of the current LLMs, for more consideration. In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al. (2019), transformer decoders use causal attention to learn causal relation signals, which is not invariant by nature Haviv et al. (2022a). Therefore, positional bias for decoder-only models is known to stem from both positional encoding and causal attention mask Yu et al. (2024); Wang et al. (2024) and is harder to mitigate.
@@ 3.1 Baseline: Order-invariant causal LMs
In this section, we briefly overview the existing work on endowing decoder-only models on order-invariance by adjusting attention mechanism, and review their limitations.
@@
A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism, by letting each segment attend to all other segments. However, to allow this within decoder-only models with causal attention and still achieve order invariance, PINE modifies the multi-head self-attention procedure to create an ‘illusion’: it treats each query segment as if it were the final segment in the input list, so that it can attend to all other segments as keys. For each query segment, the ordering among the key segments is determined by their attention scores (without positional embeddings, $\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that segments with stronger relevance to the query appear closer. Meanwhile, the tokens within a query segment still follow causal ordering. Fig. 2 illustrates this with the input [T1‘‘Given’’, S1[‘‘Apple’’], S2[‘‘Ban’’, ‘‘ana’’], S3[‘‘Orange’’], T6‘‘which one’’, .. T10‘‘red?’’]. Prefix tokens ‘‘Given’’ and suffix tokens ‘‘which one .. red?’’ remain in their original positions and follow normal causal attention, equally attending to all segments. In contrast, the segments in the middle (S1 - S3) require the order-invariant mechanism. PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID. For Case 1: When “ana” (T4, S2) acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments. Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions). Case 2: When “ana” serves as a key, the placement of S2 depends on its $\mathrm{Attn}_{\mathrm{NoPoS}}$ scores with the query segment. If “banana” (S2) is more relevant to the query segment “apple” (S1) than “orange” (S3), (i.e., $\mathbf{Attn_{NoPos}}$(S1, S2) > $\mathbf{Attn_{NoPos}}$ (S1, S3)) S2 is placed nearer to S1, with “ban” and “ana” preserving their internal sequence. For prefix, suffix, and generated tokens, they do not participate in order invariance, so they are always placed at their standard causal positions. However, the same dynamic reordering applies to segment tokens (T2–T5) when computing attention scores of suffix (T6–T10) and the generated token (T11) as query.
@@ 4 Experiment setup
Gold idx at: 0 4 9 0 4 9 14 19 0 4 9 14 19 24 29
Llama-3.1-8B-Instruct
Original 54.7 53.0 50.2 54.8 52.6 52.8 52.4 51.0 55.6 51.5 52.4 52.8 52.1 52.3 53.0
PCW 12.4 11.9 12.2 3.7 4.0 4.0 4.0 3.9 2.3 1.8 2.0 2.0 2.1 2.0 2.0
Set-Based Prompting 42.5 42.5 42.5 26.3 26.3 26.3 26.3 26.3 14.1 14.1 14.1 14.1 14.1 14.1 14.1
PINE 58.6 58.8 59.0 56.2 55.7 55.5 55.7 55.5 54.2 54.8 54.3 53.7 54.8 54.2 54.0
RoToR-lexical 61.4 61.6 61.6 61.4 59.8 59.6 59.6 59.8 59.2 59.5 59.4 59.1 59.0 59.3 59.1
RoToR-reversed lexical 61.6 61.8 61.8 58.9 59.3 58.8 58.6 58.7 57.9 58.2 57.9 57.4 57.9 57.6 57.5
RoToR-MonoT5 61.2 61.4 61.2 60.9 61.0 61.2 61.2 61.2 60.9 60.7 60.7 60.7 60.8 60.8 60.7
RoToR-Freq. 61.0 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.3 60.4 59.7 59.5 59.5 59.6 59.2
Qwen1.5-4B-Chat
Original 61.3 54.8 53.1 59.5 49.1 47.9 45.9 48.3 56.8 45.6 44.9 44.6 45.3 43.5 48.3
PINE 57.2 57.4 57.0 48.6 48.2 48.2 48.1 48.9 46.4 45.9 46.7 46.6 46.4 46.4 46.3
RoToR 58.5 58.4 58.1 49.9 49.7 49.6 49.8 49.9 44.6 44.8 44.7 44.7 44.9 44.8 44.7
RoToR-MonoT5 58.9 58.5 58.7 52.2 52.1 52.1 52.2 52.6 50.6 50.7 50.5 50.6 50.5 50.6 50.4
RoToR-Freq. 56.7 56.9 56.9 51.9 51.5 51.8 51.6 52.4 46.8 46.7 46.7 46.4 47.0 46.8 46.6
Qwen1.5-7B-Chat
Original 72.5 63.3 62.9 72.5 58.5 56.1 56.0 58.2 73.1 58.6 55.8 53.3 53.2 52.5 57.5
PINE 65.4 65.5 66.3 59.1 59.4 59.1 58.6 59.2 58.0 55.3 55.7 56.3 55.1 55.8 56.1
RoToR 68.6 68.7 68.6 62.6 62.9 62.7 63.0 62.7 57.0 57.3 59.7 57.4 57.3 62.8 57.0
RoToR-MonoT5 68.8 69.4 69.0 65.2 65.5 65.0 64.9 65.0 62.6 62.8 62.9 62.7 62.9 62.8 62.5
RoToR-Freq. 68.2 68.4 68.4 62.6 62.9 62.8 62.7 62.3 59.5 59.8 59.7 59.6 59.7 59.7 59.7
@@
Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
N = 30 N = 50 N = 30 N = 50 N = 30 N = 50
Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
Initial, no shuffling of segments
Original 50.2 44.0 51.9 50.0 44.0 51.7 30.7 27.9 34.9 31.6 28.6 35.8 31.5 27.8 35.4 31.7 28.0 35.7
PINE 51.5 45.0 52.6 51.6 45.1 52.6 31.6 28.7 35.6 31.6 28.8 35.3 32.3 28.8 36.4 32.0 28.5 35.9
RoToR 53.1 46.5 54.1 52.9 46.0 53.6 32.0 29.0 35.7 32.7 29.6 36.2 34.3 29.8 37.7 34.3 30.1 38.0
RoToR-MonoT5 51.6 45.0 52.5 52.4 45.4 52.8 32.3 29.1 36.2 32.3 29.3 35.9 32.9 28.4 36.3 32.9 28.9 36.6
RoToR-Freq. 52.6 46.1 53.7 53.1 46.4 53.7 32.3 29.2 36.0 32.3 29.2 35.9 33.7 29.5 37.2 33.5 29.5 37.2
@@
Original 49.5 43.3 51.0 49.7 43.5 51.0 30.1 27.5 34.7 30.3 27.6 35.0 31.4 27.3 35.0 31.6 27.9 35.5
$\hookrightarrow$ stdev. ($\pm$) 0.07 0.14 0.17 0.34 0.28 0.46 0.41 0.34 0.43 0.26 0.24 0.35 0.26 0.28 0.29 0.40 0.56 0.42
PINE 51.8 45.2 52.8 51.8 45.3 52.7 31.5 28.7 35.6 31.5 28.7 35.3 32.3 28.8 35.7 31.7 28.2 35.7
$\hookrightarrow$ stdev. ($\pm$) 0.05 0.07 0.16 0.15 0.16 0.19 0.20 0.18 0.13 0.17 0.20 0.21 0.17 0.20 0.13 0.18 0.16 0.14
RoToR 52.8 46.2 53.8 52.7 45.9 53.5 31.8 28.8 35.5 32.5 29.6 36.1 34.2 29.9 37.7 34.2 30.1 38.0
$\hookrightarrow$ stdev. ($\pm$) 0.05 0.05 0.02 0.05 0.09 0.04 0.05 0.02 0.09 0.11 0.06 0.09 0.09 0.07 0.06 0.06 0.05 0.04
RoToR-MonoT5 51.6 45.0 52.6 52.2 45.2 52.8 32.4 29.2 36.3 32.3 29.4 35.9 33.0 28.8 36.5 32.8 28.8 36.5
$\hookrightarrow$ stdev. ($\pm$) 0.12 0.06 0.10 0.16 0.18 0.18 0.04 0.02 0.13 0.16 0.13 0.07 0.12 0.09 0.07 0.16 0.09 0.07
RoToR-Freq. 52.5 45.9 53.5 53.1 46.4 53.7 32.3 29.3 36.0 32.4 29.3 36.1 33.8 29.6 37.4 33.7 29.6 37.4
$\hookrightarrow$ stdev. ($\pm$) 0.10 0.15 0.11 0.02 0.07 0.03 0.13 0.16 0.09 0.09 0.04 0.06 0.04 0.00 0.09 0.04 0.16 0.22
Table 2: Results on the Mintaka (KGQA) dataset on different models, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
Method Init. Rev. Avg. Init. Rev. Avg. Init. Rev. Avg.
Orig. 68.3 64.8 65.5 $\pm$ 1.0 53.6 51.9 52.6 $\pm$ 0.6 60.1 56.6 58.6 $\pm$ 0.9
@@
PINE 64.8 63.3 63.6 $\pm$ 0.7 50.5 49.3 49.4 $\pm$ 0.5 57.0 54.4 55.8 $\pm$ 0.9
RoToR 63.2 62.6 62.8 $\pm$ 0.7 49.6 47.7 48.3 $\pm$ 0.7 56.5 55.8 56.2 $\pm$ 0.6
$\hookrightarrow$ + S.R. 68.5 65.1 65.7 $\pm$ 0.9 53.7 51.8 52.6 $\pm$ 0.6 60.1 57.4 58.8 $\pm$ 0.7
RoToR - MonoT5 64.2 62.9 63.5 $\pm$ 0.5 49.7 47.6 48.7 $\pm$ 0.7 56.2 54.4 55.5 $\pm$ 0.7
$\hookrightarrow$ + S.R. 68.4 65.2 65.8 $\pm$ 0.9 53.8 51.9 52.6 $\pm$ 0.6 60.1 57.3 58.7 $\pm$ 0.8
RoToR - Freq. 64.3 63.6 63.8 $\pm$ 0.6 49.9 47.6 48.7 $\pm$ 0.5 56.4 54.7 55.7 $\pm$ 0.7
$\hookrightarrow$ + S.R. 68.5 65.3 65.8 $\pm$ 0.8 53.7 52.3 52.6 $\pm$ 0.6 60.0 57.3 58.6 $\pm$ 0.8
RoToR + S.R. (Optim.) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing (S.R), which includes both order-invariant and order-sensitive examples. RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optim.) implies the potential for further improvements.